A cannon mounted on a pirate ship fires a cannonball at 125 m/s horizontally, at a height of 17.5 m above the ocean surface. Ignore air resistance. (a) How much time elapses until it splashes into the water? (b) How far from the ship does it land?
1
Expert's answer
2011-02-01T04:45:24-0500
According to the equations of the motion
H = Ho + Voy t - gT2/2
Voy = 0;
H=0; Ho = 17.5 m; g = 10 m/s2:
T2 = 2(Ho-H)/g = 2*17.5 / 10 = 3.5
T = 1.87 s
The distance whould be
S = Vx*t = 125*1.87 = 233.85 m.